Jack is twice the age of Bill who is 8 years younger than Ben who is 4 years younger than Ted who is half as old again as Jack. How old is Bill?
Answer by ikleyn(52787)

.
Jack is twice the age of Bill who is 8 years younger than Ben who is 4 years younger than Ted who is half as old again as Jack. How old is Bill?
~~~~~~~~~~~~~~~~

Below is the line which models the condition


Jack = 2*Bill = 2*(Ben-8) = 2*((T-4)-8) = 2*(((0.5*Jack-4)-8), 


or, equivalently


Jack = Jack - 16.


The last equation is self-contradictory and does not have a solution.


Therefore, the original problem is self-contradictory and does not have a solution.